What is molality of 98% H2SO4 by wieght?

Grade:

2 Answers

Ashwin Sinha
12 years ago

Dear Akshat Sinha,

We can find the molality by the following steps:-

            Let the weight of soln. be 100 grams. then the weight of H2SO4 will be 98 grams...

            therefore,we find the no. of moles in 98% H2SO4= weight / molecular mass=98/98=1 mole

            Molality= No.of moles of solute/Wt of solvent =1/2

           Since we have taken the weight to be 100 g. We need to find it for a kg...

           =1/2*1000=500 molal soln.

 

 

 

Hope this helped you dear........

 

All The Best,.........

 

I hope that you are surely going to approve my answer by clicking 'Yes' given below..........Smile

 

SAGAR SINGH - IIT DELHI
12 years ago

 We can find the molality by the following steps:-

            Let the weight of soln. be 100 grams. then the weight of H2SO4 will be 98 grams...

            therefore,we find the no. of moles in 98% H2SO4= weight / molecular mass=98/98=1 mole

            Molality= No.of moles of solute/Wt of solvent =1/2

           Since we have taken the weight to be 100 g. We need to find it for a kg...

           =1/2*1000=500 molal soln.